# SDBA: A Stealthy and Long-Lasting Durable Backdoor Attack in Federated Learning

### v1

@@ VI Discussion
To address these issues, we investigate the primary vulnerabilities of each target model. For the LSTM model, backdoors were injected into the and layers using a layer-wise gradient masking mechanism. We empirically found that the layer is the most critical for backdoor injection, but injecting backdoors into the layer alone is insufficient to ensure durability. Due to the recurrent nature of LSTM, we also inject backdoors into the layer, which stores contextual information. Additionally, top- gradient masking is applied to enhance durability, showing that less masking improved the layer’s performance, while the layer perform best without masking. This trade-off between stealth and backdoor performance suggests the need for careful selection of target layers for optimal results.
Also, in the GPT-2 model, backdoors are primarily injected into the vulnerable layer, which has significant representational capacity for attacks. Focusing the backdoor injection on this layer proved more effective than distributing it across multiple layers, due to GPT-2’s transformer architecture, which supports parallel operations. Unlike in the LSTM model, full poisoning with only layer (without top- gradient masking) was found to be the most durable approach for GPT-2. This method reduces the likelihood of detection and ensures the backdoor persists longer, as the large-scale parameters make it less likely to be overwritten by benign clients.
@@ VII Conclusion
We proposed a novel backdoor attack method called SDBA, a stealthy and long-lasting durable backdoor attack, particularly for NLP tasks. Through a comprehensive layer-wise analysis, SDBA identified the layers most vulnerable to backdoor injection, satisfying both stealth and durability requirements. SDBA ensures the attack remains stealthy and capable of evading even six representative defense mechanisms. Moreover, by applying top-% gradient masking within the specific layers, we significantly enhance the durability of the backdoor. Our experiments demonstrated that SDBA not only surpasses existing backdoor mechanisms but also exhibits stronger durability compared to recent backdoors previously regarded as highly durable.
Future work should focus on developing robust countermeasures to combat such stealthy and durable backdoor attacks. Also, from a defense perspective, our findings reveal significant threats to FL systems.
Additionally, as mentioned in this paper, large models, such as LLMs, do not utilize distance-based metrics like those used in Multi-Krum and FLAME due to the efficiency constraints of these defense methods. In summary, there is a growing need to counter increasingly stealthy and durable attacks, alongside the challenge of developing efficient defenses suitable for large models like LLMs. This highlights the need for more powerful and efficient defenses, not only for NLP tasks but also for LLMs.

### v2

@@ VI Discussion
To address these issues, we investigate the primary vulnerabilities of each target model. For the LSTM model, backdoors were injected into the and layers using a layer-wise gradient masking mechanism. We empirically found that the layer is the most critical for backdoor injection, but injecting backdoors into the layer alone is insufficient to ensure durability. Due to the recurrent nature of LSTM, we also inject backdoors into the layer, which stores contextual information. Additionally, top- gradient masking is applied to enhance durability, showing that less masking improved the layer’s performance, while the layer performs best without masking. The results indicate that selecting appropriate target layers plays a crucial role in achieving optimal backdoor performance.
In the GPT-2 model, backdoors are injected into the layer, which functions as a key component for capturing input patterns and possesses high representational capacity. Unlike in the LSTM model, full poisoning of only the layer (without top- gradient masking) is the most durable approach for GPT-2. This method reduces the likelihood of detection and ensures that the backdoor persists longer. Furthermore, we apply the same analysis to the T5 model to evaluate how our attack is effective for other tasks such as Question Answering. The results show that our attack is more effective than existing attacks.
We also examine the additional detectability of the proposed method. While SDBA applies layer-wise gradient masking and top-% gradient masking to zero out certain gradients, this pattern alone does not constitute a practical basis for defense. Specifically, this paper assumes a scenario where the attacker intervenes after the FL training has already converged. In such a case, benign clients may also generate sparse or low-magnitude gradients due to factors such as low learning rates or non-i.i.d. local data. Consequently, detection strategies based solely on identifying sparse gradients are prone to degrading the performance of FL, undermining their reliability in practice. However, we consider the possibility of such detection, because our SDBA sets the local gradient values to exactly zero. To evade this potential detection risk while retaining the performance of backdoors, we mask the gradient using non-zero values sampled from , where is a normal distribution and denotes the standard deviation for the layer for each module containing the given coordinate.
To estimate exact statistical values, the attacker builds a surrogate model trained on their held-out benign dataset in . This surrogate model is not synchronized with the global model in FL, but it is necessary to generate statistically plausible gradients that can both evade detection and preserve the performance of SDBA.
As a result, although such defense approaches are theoretically feasible, SDBA can evade them by leveraging the surrogate model with additional computational costs. However, to the best of our knowledge, there exists no practical defense mechanism capable of detecting a single malicious client solely by comparing the global model with each local update in real-world FL scenarios. We conclude that this highlights a limitation in the generalizability of such detection strategies to various types of attacks, and also underscores the scalability challenges and computational costs associated with scanning all coordinates of local models in every training round. Therefore, in our primary attack scenario presented in Section III, we adopt zero-gradient masking as the default configuration for SDBA to reduce computational overhead while maintaining effectiveness.
@@ VII Conclusion
We proposed a novel backdoor attack method called SDBA, a stealthy and durable backdoor attack specifically designed for NLP tasks. Through a comprehensive layer-wise analysis, SDBA identifies the layers most vulnerable to backdoor injection, fulfilling both stealth and durability requirements. SDBA maintains stealthiness and is capable of evading even six representative defense mechanisms. Furthermore, by applying top-% gradient masking within specific layers, we significantly enhance the durability of the backdoor. Our experiments demonstrate that SDBA not only outperforms existing backdoor methods but also achieves greater durability than previously established attacks considered to be highly resilient.
Future work should focus on developing robust countermeasures to combat such stealthy and durable backdoor attacks. Furthermore, from a defense perspective, our findings reveal significant threats to FL systems.
Additionally, as mentioned in this paper, distance-based metrics employed in defense mechanisms such as Multi-Krum and FLAME are generally not applied to transformer-based models with a larger number of parameters than LSTM, due to the efficiency constraints of these methods. In summary, there is a growing need to counter increasingly stealthy and durable attacks, alongside the challenge of developing efficient defenses suitable for transformer-based models with significantly more parameters than RNN-based models such as LSTM. This highlights the need for more powerful and efficient defenses, not only for NLP tasks but also for transformer-based models.